If a Geiger counter is filled with 1 mol of argon gas at pressure P = 10276 Pa and temperature T = 24.4 degrees C, what is the density p of the gas in this Geiger tube in grams per cubic centimeter?


1
Expert's answer
2021-02-09T03:37:24-0500

Solution:

"ρ=\dfrac{PM}{RT}"


argon is 39.8 "(\frac{g}{mole})"


"ρ=\dfrac{(10.276)(39.8)}{(8.314)(273.15+24.4)}"


"ρ=0.165(\frac{g}{L})"


"0.165(\frac{g}{L})*\dfrac{1L}{1000cm^3}=165(\frac{g}{cm^3})"